OpenStudy (anonymous):

6) Write an equation in slope-intercept form of the line through points (-2, 6) and (7, -1). 7) Write an equation in point-slope form for the line that goes through the point (3, -8) and has a slope of -2. Can someone help i dont need anwser i just need help

OpenStudy (anonymous):

Have you drawn a quick sketch of the two points on an xy plane?

OpenStudy (anonymous):

where are you stuck?

OpenStudy (anonymous):

do u knoe formulas

OpenStudy (anonymous):

(y-y1)/(y2-y1)=(x-x1)/(x2-x1) for first que (y2-y1)=m(x2-x1)

OpenStudy (anonymous):

have u got it

OpenStudy (anonymous):

Im sorry im so confused

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1) point slope form= y-y1=m(x-x1)

OpenStudy (anonymous):

m represents slope

OpenStudy (anonymous):

m=(-1-6)/(7+2) m=-7/9 slope is decreasing now put it in point slope form can you?

OpenStudy (anonymous):

ohh okay is see how you got that

hero (hero):

(-2, 6) and (7, -1): 6 = -2m + b -1 = 7m + b b = 6 + 2m b = -7m - 1 b = b 6 + 2m = -7m - 1 7m + 2m = -6 - 1 9m = -7 m = -7/9 b = 6 + 2(-7/9) b = 6 -14/9 b = 54/9 - 14/9 b = 40/9 y = -7/9(m) + 40/9